OpenStudy (anonymous):

2x+3y=1200 as a y intercept form

pooja195 (pooja195):

Hi! Welcome to openstudy!! :D To start with we need to subtract both sides by 2x \[\huge~\rm~\bf2x-2x+3y=1200 -2x\]

pooja195 (pooja195):

@Collin2000

pooja195 (pooja195):

when you do that you are left with \[\huge~\rm~\bf 3y=1200−2x\] \[\huge~\rm~\bf 3y=-2x+1200\] Divide everything by 3 because we are solving for "Y" \[\huge~\rm~\bf 3y/3=-2x/3+1200/3\] From here you will get the slope intercept form :)

OpenStudy (anonymous):

Hint: When u want to find y-intercept\[\huge~\rm~\bf x=0\]

OpenStudy (anonymous):

Substitute\[\huge~\rm~\bf x=0\] into the equation\[\huge~\rm~\bf 2x+3y=1200\]

OpenStudy (anonymous):

\[\huge~\rm~\bf 2(0)+3y=1200\]

OpenStudy (anonymous):

\[\huge~\rm~\bf 3y=1200\]Then divide both sides by 3

OpenStudy (anonymous):

\[\huge~\rm~\bf \frac{ 3y }{ 3 }=\frac{ 1200 }{ 3 }\]

OpenStudy (anonymous):

\[\huge~\rm~\bf y=?\]solve it @Collin2000